Evaluate the cot[arccos(5-x)]
Answer by lwsshak3(11628)

Evaluate the cot[arccos(5-x)]
let A=adjacent side
let O=opposite side
H=hypotenuse
..
cot[arccos(5-x)]
This reads: cot of an angle whose cos=(5-x)
cos(5-x)=A/H=(5-x)/1
A=(5-x)
H=1
O=√(H^2-A^2) (Pythagorean Theorem)
=√[1^2-(5-x)^2]
=√[1-(25-10x+x^2)]
O=√(-24+10x-x^2)
cot[arccos(5-x)]=A/O=(5-x)/√(-24+10x-x^2)